log(s+5) = log3x + log(x+1)
find the value of x

Answer by stanbon(48535)

log(x+5) = log3x + log(x+1)
log(x+5) = log[3x(x+1)]

Take the inverse-log of both sides to get:
x+5 = 3x^2+3x
3x^2+2x-5 = 0
3x^2+5x-3x-5 = 0
x(3x+5) -(3x+5) = 0
(3x+5)(x-3) = 0
x = -5/3 or x = 3
-------
Checking:
x=-5/3; cannot be as log(3(-5/3)) = log(-5) does not exist.
x= 3: log (8) = log(9) + log(4)
log 8 = log(9*4) wrong
----------------------------
Conclusion: No solutions
===========================
Cheers,
Stan H.
===================
Cheers,
Stan H.